OpenStudy (anonymous):

Find exact value: sin(2 cos^-1(-3/11))

OpenStudy (zzr0ck3r):

there are infinite

OpenStudy (anonymous):

how so?

OpenStudy (zzr0ck3r):

cos(x) = -3/11 has infinite solutions

OpenStudy (zzr0ck3r):

forget what I said, they just want the one

OpenStudy (zzr0ck3r):

cos^(-1)(-3/11) = 105.8 now sin(105.8) = .9622

OpenStudy (zzr0ck3r):

4sqrt(7)/11

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

\[put \cos^{-1} \left( \frac{ -3 }{ 11 } \right)=x\] \[\cos x=\frac{ -3 }{ 11 }\] hence x lies in second or third quadrant. \[\sin x=\sqrt{1-\left( \frac{ -3 }{11 } \right)^{2}}=\sqrt{1-\frac{ 9 }{121 }}=\frac{ \sqrt{112} }{11 }\] sin x can be positive or negative according as x lies in second or third quadrant. plug in sin 2x= 2 sin x cos x

OpenStudy (anonymous):

wow thank you so much!

OpenStudy (anonymous):

yw

OpenStudy (zzr0ck3r):

woops I didn't notice the 2